A three-member team of University of Washington students has again been declared Outstanding Winners in the Mathematical Contest in Modeling. Of this year’s 599 participating teams, only seven, including one from the UW, were judged to be outstanding winners. In addition two other UW teams were declared Meritorious (to 10%). We have had four winning teams in the last three years.

The winning team members are Tracy Lovejoy of Kenmore, Sasha Aravkin of Bothell and Casey Schneider-Mizell of Olympia. All are seniors and double majors: Tracy and Casey in Math and Physics; Sasha, who is now a first year grad student at the UW, in Math and Computer Science. Casey is now a graduate student in Physics at the University of Michigan and Tracy is a graduate student in Physics at the UW. Last year, the UW had two Outstanding Winners and the year before, one team won. An in a way, the UW had a second win this year, since the winning team from Oxford University was advised by Jeff Giansiracusa, who was twice on winning teams as an undergrad at the UW.

The contest began at 5 p.m. on Feb. 5, when officials posted two problems on the Web. The teams had until 5 p.m. on Feb. 9 (96 hours) to select one problem and devise a solution. Competitors could access sources on the Web or in the library, but could not consult with anyone outside their team. One problem involved mathematically assessing how unique fingerprints are. The other dealt with lines at amusement parks, the problem being to find a way to issue “quick-passes” in a fair way to allow customers to avoid long waits in line.

Casey, Sasha, and Tracy chose the quick-pass problem. Their solution was an application of the idea of a Kalman filter, a concept that Sasha learned about in computer science. They gathered some of there data by observing the waiting times at Subway in the HUB. It is important for students to draw on non-standard and original ideas in the solution of the MCM problems. The team that chose to work on fingerprints used the idea of Voronoi tessellations, surely not something that would occur to most undergrads. Another UW team that worked on the quick-pass problem used ideas from economics.

All of the teams were a bit nervous since there was a tradition of winning the contest and they didn’t want to fall short of that standard. This year’s contestants might also feel the pressure, but there are many good candidates and we hope to again do quite well.

Information on the local teams, including copies of the winning papers, is available at http://www.math.washington.edu/˜ morrow/mcm/mcm.html